# https://gaoxiangjun.github.io/mani_gs/

### v1

@@ 3.4 Manipulate Gaussian Splatting through Mesh
Utilizing our triangle shape aware gaussian binding and adapting strategy, upon the completion of model training and mesh manipulation, the 3DGS is instantly manipulated and adapted. During mesh manipulation, the attributes in the local triangle space remain unchanged. The triangle rotation, position, and edge length can be calculated instantly. Therefore, the global Gaussian position, scaling, and rotation can be self-adaptively adjusted following our proposed formula. In this paper, we exhibit the 3DGS manipulation rendering outcomes, such as large-scale deformation, local manipulation, and soft-body simulation, which are driven by the manipulated mesh. Many 3D design software applications possess the capability to execute mesh manipulation. In our experiments, we employ Blender to manipulate the mesh.
@@ 4.1 Training Details
The first stage of our methods includes a mesh extracting stage, during which we extract triangular mesh from neural implicit surface filed (NeuS[31]) or 3D Gaussian splatting (Screened Poisson or Marching Cube) only. However, the extracted mesh always contains enormous triangles, which we try to decimate to approximately 300K.
With the extracted mesh, we conduct the triangle shape aware Gaussian binding and adapting strategy on the mesh. For each triangle, we bind $N=3$ Gaussian on the surface initially. The Gaussian attributes are optimized subsequently with the supervision of multi-view rendering loss in the second stage. We train our model for 30K iterations in the initial stage to extract mesh and 20K iterations in the second stage. All experiments are conducted on a single NVIDIA A100 GPU. Please see the supplementary material for more details.
@@ 4.2 Datasets, Metrics and Methods for Comparison
To evaluate our methods, we compare Mani-GS with previous editable novel view synthesis methods, a NeRF-based editing method NeRF-Editing [19] and a 3DGS-based editing method SuGaR[7]. For the evaluation, we employ the commonly used metrics:PSNR, SSIM, LPIPS. We evaluate our methods mainly on the NeRF Synthetic dataset[21] and DTU dataset[11].
@@ 4.3 Evaluation
Static Rendering Table 1 provides a quantitative comparison of all NeRF Synthetic 8 cases between our method and competing methods. We conducted experiments using their official code repository.
Table 1: Quantitative comparison of our methods with NeRF-Editing (N.E.)[19] and SuGaR[7] on NeRF Synthetic dataset in terms of SSIM, PSNR, LPIPS. ($\uparrow$ means higher is better, $\downarrow$ means lower is better.) The best results are marked in bold.
@@
The numerical results of the NeRF-Editing were not presented in their original paper, we ran their code to provide more visual and numerical comparisons. As depicted in Table 1, we observed some outliers in "Drums, Ficus" that were lower than even 10 PSNR compared to ours and SuGaR. Therefore, we remove these outliers with a strikeout in the table. As can be observed, our approach surpasses all the baseline methods with respect to PSNR, SSIM, LPIPS, which means we achieve the best rendering quality. We are 3.0 higher than SuGaR in PSNR, 0.013 higher in SSIM, and 0.024 lower in LPIPS.
Figure 3: Visual comparison between ours, NeRF-Editing[19](N.E.) and SuGaR[7] for static rendering. It illustrates our proposed method can contain a much more accurate boundary in "Ship" (first row), and detailed results in "Drums" (second row).
In Figure 3, we present qualitative results of our approach and other methods in overview and zoom-in details. Our methods render more detailed and more accurate boundaries. For SuGaR, it attempts to bind 3D Gaussians on the triangle and enforce that the attached Gaussian is closely aligned on the triangle. In practice, they set the scale $s_{1}$ along the triangle’s normal direction to a value close to zero, i.e., $s_{1}=\epsilon$. This binding strategy heavily depends on the accuracy of the mesh. As can be observed in the third column of Figure 3, wrong geometry leads to an inaccurate rendering, especially in the boundary region.
Figure 4: We provide a visual comparison of manipulation rendering between our proposed approach and SuGaR. Our approach exhibits fewer artifacts and less blurring effect compared to SuGaR. For more details, please zoom in.
Manipulation Rendering In Figure 4, we showcase our manipulation results. In these four cases, we manipulate the underlying mesh with large deformation, the Chair is stretched, Lego is tapered, Ficus and Mic is bent to left and up respectively. As demonstrated in Chair, Ficus, we have more accurate boundary and shape, as well as in the bottom region. This indicates that when the geometry is not that accurate, SuGaR can not adapt to compensate for geometry error, which results in missing geometry and dilated boundary geometry. For Lego, Mic, we can maintain the high rendering quality even after the large deformation, while SuGaR shows some distortion and noise in rendering results after large deformation. We did not present the manipulation results of NeRF-Editing[19] since we did not obtain their relatively reasonable results in these cases.
Figure 5: Visual results of local manipulation and soft body simulation. In row 1, we demonstrate that our proposed GS-mesh binding method enables us to support local part manipulation, such as blending, reposing, and elastic deformation. In row 2, we showcase a 3DGS soft body simulation demo at different timesteps.
In the second row of Figure 5, we present the rendering results of soft body simulation at different timesteps. As observed, we can achieve 3DGS soft body simulation by just transferring the mesh simulation to 3DGS, which eliminates the need for a soft body simulation algorithm dedicated to 3DGS.
Table 2: Quantitative ablation comparison between 3DGS On NeuS Mesh, NeuS Mesh + Offset, Ours with Marching Cube Mesh,Ours with Screened Poisson Mesh on NeRF Synthetic dataset. ($\uparrow$ means higher is better, $\downarrow$ means lower is better.)
[…]
3DGS On NeuS Mesh 30.87 0.9521 0.0447
NeuS Mesh + Offset 32.48 0.9625 0.0341
@@
4.4 Ablation Study
We conduct ablation studies to verify the effectiveness of triangle shape aware Gaussian binding and adapting method. We first evaluate the strategy of directly binding 3DGS to the mesh, which implies that the 3D position is fixed on the triangle. As shown in Table 2, the performance significantly drops, with a decrease of approximately 2.6 PSNR compared to our best model. For visual ablation in Figure 6, 3DGS on NeuS Mesh after deformation shows a boundary with many burrs.
Next, we verify the effectiveness of adding 3D offset for 3DGS on Mesh. Although the offset can enhance the fitting of 3DGS to the static scene, as demonstrated in Table 2, it fails to generate satisfactory deformation rendering results because the offset only fits the static scene and remains unchanged during subsequent deformations. Consequently, it leads to significant noise and distortion after manipulation in Table 2.
Figure 6: Visual Results of Ablation Study. After deformation, (3DGS on Mesh) shows a burring boundary, (Mesh + Offset) leads to significant noise and distortion, (Ours Full) can maintain the high fidelity rendering. In the second row, we demonstrate that even with a low-quality mesh, we can still achieve high-quality editable rendering.
Finally, we conduct experiments with different meshes extracted using different methods. The 3DGS Marching Cube Mesh (3DGS MC) is of low quality, including a dilated boundary and very noisy surface, as can be observed in Figure 6 row 2. The screened poisson mesh (Poisson Recon.) has some unconnected regions and missing parts compared with NeuS mesh (NeuS). However, using our triangle shape aware Gaussian binding and adapting method can still achieve 3DGS manipulation and maintain high-fidelity rendering even after very large deformation as shown in Figure 6 row 2. As shown in Table 2, the numerical results obtained with screened poisson mesh are only slightly lower than those obtained with NeuS mesh. When the mesh is of low quality, such as the MC mesh, the quantitative results are approximately 1 PSNR lower than the best, but still 2 PSNR higher than only binding 3DGS on the best Mesh (NeuS mesh), and 1.5 PSNR higher than SuGaR.
We also evaluate our method using mesh extracted from SuGaR, whose mesh is extremely dense. The PSNR using SuGaR mesh is 33.67, which is about 3 PSNR higher than that of the original SuGaR results.
5 Conclusion and Limitation
In this paper, we introduce a triangle shape aware Gaussian binding strategy with self-adaptation, which supports various 3DGS manipulations, maintains rendering quality, and has a high tolerance for mesh accuracy. We evaluate our methods on the NeRF synthetic dataset and demonstrate state-of-the-art results, showcasing various 3DGS manipulations, including large deformations, local manipulations, and soft body simulations.
During our experiments, we noticed that some results still exhibit distortions. When the local region of the manipulated mesh contains highly non-rigid deformations, it can result in rendering distortions. Additionally, during our simulation demos, we found that conducting physics simulations on meshes with more than 35K triangles can take hours. It would be a novel direction to explore fast simulation methods for 3DGS. Finally, we found that our results may not have accurate boundary rendering when the extracted mesh has a significant discrepancy from the ground truth mesh, such as unconnected regions.
@@ 5 Conclusion and Limitation
In this supplementary material, we describe more implementation details of our method in Sec. A. Then, we give an efficiency analysis of training time and inference time in Sec. B. Besides, we also provide more visual rendering and geometry results on the NeRF[21] Synthetic dataset including some video demos in Sec. C. Finally, we further evaluate our methods on the DTU[11] dataset and provide qualitative and quantitative results in Sec. D.
Appendix A Implementation Details
A.1 Training Details of Mesh Extraction Stage
@@
In this work, we try to extract triangular mesh using the Screened Poisson surface reconstruction [13] method from trained Gaussian-Splatting model. We incorporate a normal attribute $\bm{n}$ for each 3D Gaussian and optimize the normal attribute with the pseudo-normal constraint.
@@
A.2 Training Details of Gaussian-Binding Stage
@@
In the Gaussian-Binding stage, we don’t perform adaptive control because we find it doesn’t influence the final performance. We also train 300K iterations in this stage with L1 loss, SSIM loss and mask entropy loss. The overall loss in this can be summarized as follows:
@@
Appendix B Efficiency Analysis
The efficiency of GS Binding training and rendering speed depends on the number of Gaussians, which is the product of the triangle number $T$ and the Gaussians number for each triangle $N$. In Table C, We first fixed $T$ and tested different values of $N$. Our results indicate that $N$=3 leads to the best rendering quality while keeping a competitive rendering speed. When $N=1$, the PSNR slightly decreased with a faster training and rendering speed.
We also evaluated the impact of underlying mesh resolution by testing meshes with different triangles (270K, 150K, 70K). As shown in Table C, the rendering quality decreases while efficiency improves with decreasing mesh resolution.
Table C: Efficiency Analysis
$N$=4 $N$=3 $N$=1 $N$=1 $N$=1
Triangles(K) 270 270 270 150 70
Points(K) 1080 810 270 150 70
Training (min) 16 13 7 5.5 4.5
Speed (FPS) 244 300 452 571 572
PSNR 36.36 36.39 36.27 35.86 34.52
Appendix C More Results on NeRF Synthetic Dataset
. Soft Body Simulation. In addition to the visual results presented in Figure 5 of the main paper, we also provide the geometry after simulation and rendering at different viewpoints in Figure G. To improve the speed of the mesh simulation, we decimated the original mesh from 300K to 35K triangles. While this may result in some decrease in rendering quality due to the reduced number of triangles as well as Gaussians, it was necessary to ensure reasonable simulation speed.
Demo Video. In order to further demonstrate the effectiveness of our methods, we have provided additional visual videos showcasing large deformation, soft body simulation, and local manipulation. These videos can be accessed through a local webpage by navigating to the mani-gs folder and clicking on index.html.
Numerical Comparison with NeuS and Neu-Mesh. We have compared our methods with NeuS and Neu-Mesh on the NeRF Synthetic dataset, and the results are presented in Table D. Our approach achieves a PSNR score that is 2.5 higher than Neu-Mesh, and a SSIM score that is 0.014 higher than Neu-Mesh, and a LPIPS score that is 0.012 lower than Neu-Mesh. These results demonstrate that our approach has achieved the best overall performance in all metrics compared to NeuS and Neu-Mesh on the NeRF Synthetic dataset.
Binding Gaussians on low-quality mesh. In some extreme cases, the Screend Poisson surface reconstruction method may result in a very low-quality triangular mesh. However, with our mesh-Gaussians binding strategy, we can still generate much better rendered images than those produced by SuGaR [7], even though SuGaR may have a much better mesh in such case as shown in Figure H.
Figure G: Visual results of softbody simulation at different viewpoints. The left column displays the geometry after simulation, while the right three columns showcase the rendering results from three different viewpoints.
Figure H: Binding Gaussians on a low-quality mesh (ours on the left, SuGaR on the right), we are still able to achieve high-fidelity manipulated rendering results when the mesh we generate of Screened Poisson reconstruction is of low quality. In contrast, SuGaR fails to produce satisfactory results, even though it has a better mesh in this particular case.
Appendix D More Results on DTU Dataset
Methods DTU NeRF 360 Synthetic
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
NeuS [31] 26.352 0.909 0.176 30.588 0.960 0.058
Neu-Mesh [38] 28.289 0.921 0.117 30.95 0.951 0.043
Ours 31.496 0.943 0.088 33.45 0.965 0.031
Table D: We compare quantitative rendering quality with NeuS [31] and NeuMesh [38] on the DTU dataset and the NeRF Synthetic dataset.
We have also evaluated our approach on the DTU dataset, which includes 15 cases with multi-view images as input. The results are presented in Table D, and demonstrate that our approach achieves a PSNR score that is 2.8 higher than Neu-Mesh, and a SSIM score that is 0.022 higher than Neu-Mesh. Additionally, our approach has a LPIPS score that is 0.016 lower than Neu-Mesh, indicating that our approach achieves the best overall performance on the DTU dataset.
We have also presented the manipulation results in Figure I. The left two columns showcase the geometry and rendered image before manipulation, while the right three columns showcase the geometry and rendered image after manipulation. In the first row, the wing of a pigeon is manipulated, while in the second row, the arms of a tiger are swinging down. In the third row, a toy is twisted to the left. The manipulation results presented in Figure I demonstrate that our approach can successfully transfer mesh manipulation to Gaussian-Splatting, resulting in accurate and visually appealing results.
Figure I: Manipulation rendering results in DTU dataset. The left two columns showcase the geometry and rendered image before manipulation, while the right three columns showcase the geometry and rendered image after manipulation. To highlight the deformed area, we have enclosed it within a red rectangle.

### v2

@@ 3.4 Manipulate Gaussian Splatting through Mesh
Utilizing our triangle shape aware gaussian binding and adapting strategy, upon the completion of model training and mesh manipulation, the 3DGS is instantly manipulated and adapted. During mesh manipulation, the attributes in the local triangle space remain unchanged. The triangle rotation, position, and edge length can be calculated instantly. Therefore, the global Gaussian position, scaling, and rotation can be self-adaptively adjusted following our proposed formula. In this paper, we exhibit the 3DGS manipulation rendering outcomes, such as large-scale deformation, local manipulation, and soft-body simulation, which are driven by the manipulated mesh. In our experiments, we employ Blender to manipulate the mesh.
@@ 4.1 Training Details
Figure 4: We offer an editing comparison between our method, SuGaR, and NeuMesh. Our approach demonstrates fewer artifacts and less blurring effects than SuGaR, and presents more abundant and distinct details compared to NeuMesh. For further details, please zoom in.
The first stage of our methods includes a mesh extracting stage, during which we extract triangular mesh from NeuS [37] or 3DGS (Screened Poisson or Marching Cube). However, the extracted mesh always contains enormous triangles, which we try to decimate to around 300K.
With the extracted mesh, we conduct the triangle shape aware Gaussian binding and adapting strategy on the mesh. For each triangle, we bind $N=3$ Gaussian on the surface initially. The Gaussian attributes are optimized subsequently with the supervision of multi-view rendering loss in the second stage. We train our model for 30K iterations in the initial stage to extract mesh and 20K iterations in the second stage. All experiments are conducted on a single NVIDIA A100 GPU.
@@ 4.2 Datasets, Metrics and Methods for Comparison
To evaluate our methods, we compare Mani-GS with previous editable novel view synthesis methods, NeRF-based editing method NeRF-Editing  [51], NeuMesh  [45] and 3DGS-based editing method SuGaR [10]. For the evaluation, we employ the commonly used metrics:PSNR, SSIM, LPIPS. We evaluate our methods mainly on the NeRF Synthetic dataset [27] and DTU dataset [16].
@@ 4.3 Evaluation
Static Rendering Table 1 provides a quantitative comparison of all NeRF Synthetic 8 cases between our method and competing methods. We conducted experiments using their official code repository. We remove two outliers (“Drums, Ficus”) from NeRF-Editing using a strikeout. As observed, our approach outperforms all baseline methods in terms of PSNR, SSIM, and LPIPS, indicating that we achieve the best rendering quality. Our method achieves SOTA results across all 8 cases, surpassing SuGaR by at least 1.5 points in PSNR when using NeuS mesh. For a fair comparison with SuGaR, we also evaluate our binding method using the same mesh as SuGaR (denoted as “Ours(S*)”). Our method maintains its superior performance even with SuGaR mesh, showing comparable results to those with NeuS mesh.
In Figure 3, we present qualitative results of our approach and other methods in overview and zoom-in details. For SuGaR, it attempts to bind 3D Gaussians on the triangle and enforce that the attached 3DGS is as flat as possible and is closely aligned with the triangle. This binding strategy heavily depends on the accuracy of the mesh. As can be observed in the third column of Figure 3, wrong geometry leads to an inaccurate rendering, especially in the boundary region. N.E.[25] results are blurred and lack intrinsic details. In contrast, our method achieves higher fidelity rendering with more precise boundaries and richer details.
Methods DTU
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
NeuMesh [45] 28.29 0.921 0.117
SuGaR [10] 30.06 0.921 0.136
Ours 31.50 0.943 0.088
Table 2: We compare quantitative rendering quality with SuGaR and NeuMesh on the DTU dataset. Note that ours and SuGaR employ the same underlying mesh extracted from poisson recon.
We also evaluate our methods on the DTU dataset [16] with real objects, as shown in Table 2. It is important to note that we use the same base mesh as SuGaR in this dataset. As listed, we outperform both SuGaR and NeuMesh across all metrics, further demonstrating the efficacy of our 3DGS binding strategy.
Figure 5: Visual comparison of inaccurate mesh binding and rendering. Note that ours and SuGaR employ the same underlying mesh extracted from poisson recon.
Binding 3DGS on inaccurate mesh. When the mesh accuracy is low, as shown on the left side of Fig 5, SuGaR produces distorted results due to the missing geometry. Conversely, when the mesh accuracy is high but still deviates from the ground truth mesh, SuGaR tends to generate blurred results because of view inconsistency with the given mesh. Thanks to our binding strategy in the triangle’s local space, which allows 3DGS to move freely around the triangle, we can achieve high-fidelity rendering in both scenarios. This demonstrates that our method exhibits a high tolerance for mesh accuracy.
Figure 6: The manipulated geometry and corresponding rendering results of our methods on DTU dataset.
Manipulation Rendering In Figure 4, we showcase our manipulation results. In these four cases, we manipulate the underlying mesh with large deformation, the Chair is stretched, Lego is tapered, Ficus and Mic is bent respectively. As demonstrated in Ficus and Mic, we have more accurate boundaries and shapes, as well as distinct details. However, SuGaR can not adapt to compensate for geometry errors, which results in missing or dilated boundary rendering with distortions. For Lego and Chair, we can maintain the high rendering quality even after the large deformation, while NeuMesh showcases blurred and noised results.
Figure 7: Visual results of local manipulation (row 1) and soft body simulation at different timesteps.
In row 2 of Figure 7, we present the rendering results of soft body simulation at different timesteps. As observed, we can achieve soft body simulation by just transferring the mesh simulation to 3DGS, which eliminates the need for a soft body simulation algorithm dedicated to 3DGS.
4.4 Ablation Study
Figure 8: Ablation of binding strategy and geometry. After deformation, (3DGS on Mesh) shows a burring boundary, (Mesh + Offset) leads to significant noise and distortion, (Ours Full) can maintain the high fidelity rendering. In the second row, we demonstrate that even with a low-quality mesh, we can still achieve high-quality editable rendering.
We conduct ablation studies to verify the effectiveness of triangle shape-aware Gaussian binding and adapting method. We first evaluate the strategy of directly binding 3DGS on mesh, which implies that the 3D position is fixed on the triangle. The mesh is extracted from NeuS. As shown in Table 3, the performance significantly drops, with a decrease of approximately 2.6 PSNR compared to our best model. For visual ablation in Figure 8, 3DGS on Mesh after deformation shows a boundary with many burrs.
Next, we verify the effectiveness of adding 3D offset for 3DGS on Mesh (Mesh + Offset). Although the offset can enhance the fitting of 3DGS to the static scene, as demonstrated in Table 3, it fails to generate satisfactory deformation rendering results because the offset only fits the static scene and remains unchanged during subsequent deformations. Consequently, it leads to significant noise and distortion after manipulation in Figure 8, whereas our full model can maintain high-quality rendering after manipulation, demonstrating the efficacy of the 3DGS binding strategy in local triangle space.
Table 3: Quantitative ablation comparison between 3DGS On Mesh, Mesh + Offset, Ours with Marching Cube Mesh or Screened Poisson Mesh on NeRF Synthetic dataset. ($\uparrow$ means higher is better, $\downarrow$ means lower is better.)
@@ 4.4 Ablation Study
3DGS On Mesh 30.87 0.9521 0.0447
Mesh + Offset 32.48 0.9625 0.0341
@@
Finally, we conduct experiments with different meshes extracted using different methods. As can be observed in Figure 8 row 2, 3DGS Marching Cube Mesh (3DGS MC) is of low quality, including a dilated boundary and very noisy surface. And the screened poisson mesh (Poisson Recon.) has some unconnected regions and missing parts compared with NeuS mesh. By employing our triangle shape aware Gaussian binding and adapting method, we can still achieve 3DGS manipulation and maintain high-fidelity rendering after large deformation. In Table 3, the numerical results obtained with screened poisson mesh are only slightly lower than those obtained with NeuS mesh. When the mesh is of low quality (Marching Cube mesh), the quantitative results are approximately 1.3 PSNR lower than the best, but still 1.3 PSNR higher than only binding 3DGS inside the triangle of the best Mesh.
5 Conclusion
In this paper, we introduce a triangle shape aware Gaussian binding strategy with self-adaptation, which supports various 3DGS manipulations, maintains rendering quality, and has a high tolerance for mesh accuracy. We evaluate our methods on both synthetic and real datasets and demonstrate SOTA results.
During our experiments, we noticed that some results still exhibit distortions. When the local region of the manipulated mesh contains highly non-rigid deformations, it can result in render distortions. Additionally, during our simulation demos, we found that conducting physics simulations on meshes with more than 35K triangles can take hours.
@@ 5 Conclusion
Overview. The supplementary material has the following contents:
  • More visual results
  • Efficiency Analysis
  • Implementation Details
A. More Visual Results
Demo Video. In order to further demonstrate the effectiveness of our methods, we have provided additional visual videos showcasing large deformation, soft body simulation, and local manipulation. These videos can be accessed through project page.
Soft Body Simulation. In addition to the visual results presented in the main paper, we also provide the geometry after simulation and rendering at different viewpoints in Figure 9. To improve the speed of the mesh simulation, we decimated the original mesh from 300K to 35K triangles. While this may result in some decrease in rendering quality due to the reduced number of triangles as well as Gaussians, it was necessary to ensure reasonable simulation speed.
Figure 9: Visual results of softbody simulation at different viewpoints. The left column displays the geometry after simulation, while the right three columns showcase the rendering results from three different viewpoints.
More real case. We have also presented more manipulation results of real scenes in Figure 10. The top three images are from the DTU dataset, while the bottom two are from the Tanks and Templates dataset[21]. Images within the dotted black rectangles are before manipulation, whereas those marked with red arrows or rectangles are after manipulation. The manipulation results presented in Figure 10 demonstrate that our approach can successfully transfer mesh manipulation to Gaussian-Splatting, resulting in accurate and visually appealing results.
Figure 10: Manipulation rendering results on real object dataset. To highlight the deformed area, we have enclosed it within a red rectangle.
Table 4: Efficiency Analysis
$N$=4 $N$=3 $N$=1 $N$=1 $N$=1
Triangles(K) 270 270 270 150 70
Points(K) 1080 810 270 150 70
Training (min) 16 13 7 5.5 4.5
Speed (FPS) 244 300 452 571 572
PSNR 36.36 36.39 36.27 35.86 34.52
B. Efficiency Analysis
The efficiency of 3DGS Binding training and rendering speed depends on the number of Gaussians, which is the product of the triangle number $T$ and the Gaussians number for each triangle $N$. In Table 4, We first fixed $T$ and tested different values of $N$. Our results indicate that $N$=3 leads to the best rendering quality while keeping a competitive rendering speed. When $N=1$, the PSNR slightly decreased with a faster training and rendering speed.
We also evaluated the impact of underlying mesh resolution by testing meshes with different triangles (270K, 150K, 70K). As shown in Table 4, the rendering quality decreases while efficiency improves with decreasing mesh resolution.
C. Implementation Details
C.1 Training Details of Mesh Extraction Stage
@@
In this work, we try to extract triangular mesh using the Screened Poisson surface reconstruction [18] method from a trained Gaussian-Splatting model. We incorporate a normal attribute $\bm{n}$ for each 3D Gaussian and optimize the normal attribute with the pseudo-normal constraint.
@@
C.2 Training Details of Gaussian-Binding Stage
@@
In the Gaussian-Binding stage, we don’t perform adaptive control because we find it doesn’t influence the final performance. We also train 300K iterations in this stage with L1 loss, SSIM loss, and mask entropy loss. The overall loss in this can be summarized as follows: